A portfolio is invested 20 percent in Stock G, 27 percent in Stock J, with remainder in Stock K. The expected returns on these stocks are 9.45 percent, 13.83 percent, and 17.89 percent, respectively. What is the portfolio's expected return?  Answer to two decimals. Thank you!!
Answer #1

remainder in Stock K=100-(20+27)=53%

Expected return=Respective return*Respective weight

=(0.2*9.45)+(0.27*13.83)+(0.53*17.89)

=15.11%(Approx)